Calculate the frequency and energy of a photon that has a wavelength of 1075.0nm.
Answer #1

We have

wavelength \lambda = 1075.0 nm = 1075 \times 10-9 m

We know the formula

frequency (v) = Speed of light (c) / wavelength

where speed of light = 3.00 \times 108 m/s

So by solving with given values we get

frequency (v) = 3.00 \times 108 m/s / 1075 \times 10-9 m = 2.78 \times 1014 Hz

Now for energy we have formula

Energy (E) = hc/ \lambda

Where h = Plancks constant = 6.626 \times 10- 34 J.s

So we get

Energy = 6.626 \times 10- 34 J.s \times  3.00 \times 108 m/s / 1075 \times 10-9 m

= 1.85 \times 10- 19 J